Find the term not involving y in the expansion of
bracket( 1 over y exponent 3 + y exponent 2 closed bracket) exponent 5

Answer by stanbon(48546)

Find the term not involving y in the expansion of
bracket( 1 over y exponent 3 + y exponent 2 closed bracket) exponent 5
------------------
[(1/y^3) + y^2]^5
= [[y^5+1]^5]/y^15
Let m = y^5
= [m+1]^5/m^3
= [m^5 + 5m^4 + 10m^3+.....]/m^3
Comment: Here you can see that if you divide thru by m^3 you will
have a term 10m^3/m^3 = 10
----------------------------
That is the term that has no "m" in it and therefore no "y".
That term is 10m^3/m^3 = 10y^15/y^15
============================
Cheers,
Stan H.